OpenStudy (anonymous):

One leg of a triangle has a measure of 24 units. If the hypotenuse measures 25 units, what is the length of the other leg?

OpenStudy (johnweldon1993):

 remember the pythagorean theorem? It states that \[\large a^2 + b^2 = c^2\] for right triangles...where c is the length of the hypotenuse...and 'a' and 'b' are the lengths of the other legs of the triangle... so since we know the hypotenuse 'c' and one leg 'a' we can solve for 'b' *I've labeled as 'x' here sorry* \[\large a^2 + x^2 = c^2\] How can we rearrange that to solve for 'x' ?

OpenStudy (anonymous):

Use the Pythagorean Theorem.